I must decide if the algebraic manipulation is correct or incorrect. if it is incorrect I need to fix it or explaine why it cannot be simplified.I don't know if this is the right spot to post but if you can help that would be great.
6x^2-14x+4=(3x-4)(2x-1)

Answer by Alan3354(21592)

I must decide if the algebraic manipulation is correct or incorrect. if it is incorrect I need to fix it or explain why it cannot be simplified.I don't know if this is the right spot to post but if you can help that would be great.
6x^2-14x+4=(3x-4)(2x-1)
------------------
That's not correct.
6x^2-14x+4 = 3x^2 - 7x + 2 (divide by 2)
= (3x-1)*(x-2)
Or, for the original, double either binomial:
6x^2-14x+4 = (6x-2)*(x-2), or
6x^2-14x+4 = (3x-1)*(2x-4)
--------------------
(3x-4)*(2x-1) = 6x^2 - 11x + 4, not the same.